Fill in the blanks:


(i) One metre is ______________ cm.


(ii) Five kilometre is ______________ m.


(iii)Motion of a child on a swing is ______________.


(iv) Motion of the needle of a sewing machine is _____________.


(v) Motion of wheel of a bicycle is______________.


Grade:upto college level

11 Answers

Navjyot Kalra
10 years ago

(i) One metre is __100__ cm.

(ii) Five kilometres is __5000__ m.

1 km = 1000 m

5 km = 1000 × 5 = 5000 m

Hence, the answer is 5000 m.

(iii) Motion of a child on a swing is __periodic__.

KARTHIK
10 years ago
(i) One metre is=100 cm. (ii) Five kilometre is =5,000 m.
Adarsh
8 years ago
  1. 100 cm
  2. 5000m
  3. periodic
  4. periodic
  5. rotational
Prakshi Singh
7 years ago
1m = 100 cm
2] 5 km is 5000m
1 km = 1000m
5 km  =  1000*5= 5000m
hence, the answer is 5000 m
motion of a child swing is periodic.
motion of a needle of sewing machine is periodic.
motion of a wheel of a bicycle is rotational.
Ritika Das
6 years ago
hundred centimetre
five thousand metre
simple harmonic or periodic
periodic
oscillatory or rotational
Amarjit
6 years ago
The answer to the following questions are as follows 1) 1002) 50003) periodic4) periodic5) circular
Nitesh
6 years ago
100 cm5000 mPeriodic motion !! ! !! ! ! ! ! ! ! !! ! ! ! ! !!! !! !! ! !! ! ! ! Aa a a. A aa a a a. A aa a. A a a a a a. A a a. A. A. A
Harshit Panwar
6 years ago
1m = 100cm
5km = 5000m
Periodic motion
Periodic motion
Circular or oscilatory or rotational motion
Hope it helps
Kochouseph
6 years ago
Answers for your questions (I) 100 centimeters(Ii)5000meters(III) periodical movement(IV) periodical movement(V) rotational movement
praveen kumar
6 years ago
1 meter = 100 cm
5 km =5*1000=5000m
forced oscillation.
simple harmonic motion.
circular and translatory motions.
Priyansh
6 years ago
100 cms
5000m
Periodic
Periodic
Rotational.